OpenStudy (joannablackwelder):

When a sample of moon rock was analyzed by mass spectroscopy, the ratio of K-40 to Ar-40 was found to be 0.1295. Based on this ratio, how old is the moon?

OpenStudy (aaronq):

It's called "mass spectrometry" :P but knowing that the decay is a first-order process, we can use: \(\sf \large A_t=A_o*e^{-kt}\) Where \(A_t\) is the amount at time \(t\) \(A_o\) is the initial amount k is the decay constant t is time You'll need to find k first using the half-life, \(t_{1/2}\): \(t_{1/2}=\dfrac{ln(2)}{k}\)

OpenStudy (aaronq):

The ratio they're talking about is the rearragement of \(A_t\) and \(A_o\) \(\sf \large A_t=A_o*e^{-kt}\rightarrow \dfrac{A_t}{A_o}=e^{-kt}\)

OpenStudy (aaronq):

wait i think the ratio is actually the reverse

OpenStudy (joannablackwelder):

Right, but how does that relate to the ratio of K-40 and Ar-40?

OpenStudy (joannablackwelder):

I found that K-40 only decomposes to Ar-40 part of the time.

OpenStudy (joannablackwelder):

https://en.wikipedia.org/wiki/Potassium-40

OpenStudy (aaronq):

so it's 10.72% of the time, you'll have to account for that and the half-life is 1.251(3)×10^9 years let me think about this for a min

OpenStudy (joannablackwelder):

Sure :-)

OpenStudy (aaronq):

Is the ratio \(\dfrac{[Ar]}{[K]}=\dfrac{1}{0.1295}\) or \(\dfrac{[Ar]}{[K]}=\dfrac{0.1295}{1}\) ?

OpenStudy (joannablackwelder):

Your guess is as good as mine. The problem reads just as I have written it.

OpenStudy (aaronq):

I tried it both ways, this way gives me a positive value the other gives me a negative, so \(\dfrac{[Ar]}{[K]}=\dfrac{1}{0.1295}=e^{-kt}\) need to take account of all the products: \(\dfrac{[Ar]}{0.1072}=\dfrac{[All~products ]}{1}\rightarrow [Ar]=\dfrac{[All]*0.1072}{1}\) Back into original equation: \(\dfrac{[Ar]}{[K]}=\dfrac{[All]*0.1072}{[K]}=\dfrac{0.1072 }{0.1295}\) So \(\huge \dfrac{0.1072 }{0.1295}=e^{-0.0000000001846915t}\) t=1,023,244,884 years 1 billion years sounds plausible

OpenStudy (aaronq):

i have to go, i'm not 100% on this, i can take a look at it later if you're not satisfied

OpenStudy (joannablackwelder):

That make a lot of sense at first glance. Thanks so much!

OpenStudy (aaronq):

no problem !